Skip to end of metadata
Go to start of metadata

You are viewing an old version of this page. View the current version.

Compare with Current View Page History

Version 1 Current »

Summary:  LICORs at V06 and V07 still going down. PIO is being used to restart them. Need to verify electronic boxes are grounded.
                  TRH.20m on RSW06 down.
                  Tsoil at TNW05 down.
                   RNE04 tower:  Discussions took place yesterday about the safety of the tower. The latest plan is to have the questionable anchor replaced and the guywires tighten to a minimum of 400. Hopefully this will happen today. Steve and I went to the site yesterday to install a power line, the DSM, and the WIFI link. Power came from a LIDAR approximately 85m from the tower. As soon as the tower is made safe the remaining hardware will be installed.

 

 

  • No labels